# A Survey of AI Scientists

### v4

@@ 4.4 Physics and Engineering
    SR-Scientist [SRScientist2025] employs agentic workflows for symbolic regression, autonomously discovering fundamental equations from observational data.
@@
    AI Feynman [Udrescu2020AIPhysRev] introduces physics-constrained symbolic search (dimensional consistency), rediscovering classical laws from raw data and inspiring modern symbolic reasoning architectures.
@@
    Quantum-Agent-SDL [Cao2024QuantumAgentSDL] combines reinforcement learning with LLM-based hypothesis refinement for self-optimized qubit calibration and error correction in quantum computing.
@@ 4.5 Meta-Science and Social Science
    SciAgents [Ghafarollahi2024SciAgents] employs multi-agent collaboration and dynamic graph reasoning to traverse publication/author networks, identifying knowledge gaps, interdisciplinary connections, and predicting research trends.
@@
    AI for Social Science (AI4SS) [xu2024ai] outlines roadmaps for AI-driven large-scale social modeling, policy simulation, and analysis of innovation diffusion and collaboration patterns.
@@
    Ethical Governance Frameworks [resnik2025ethics, Mwita2025AIWritingEthics] address societal implications of AI-generated science, proposing frameworks for authorship management, credit attribution, and accountability maintenance.
@@ 5 Open Problems and Future Directions
  • Reproducibility-by-Design and Verifiable Science. Multi-stage autonomous systems amplify reproducibility challenges through compositional complexity, where parameter variations in early stages cascade into divergent outcomes. Achieving verifiable science requires architectural integration of three components: (1) Environmental Determinism—containerization of computational dependencies, cryptographic hashing of data artifacts, and version control of model checkpoints ensure bitwise reproducibility; (2) Fine-Grained Provenance—cryptographic linking of manuscript claims to specific code commits, data versions, and model parameters enables claim-level verification [Gu2024BLADE]; and (3) Automated Verification—AI-driven auditor agents performing logical consistency checks, detecting claim-evidence misalignments, and validating statistical inferences before publication [son2025ai]. These mechanisms must transition from post-hoc validation to intrinsic architectural properties.
@@
  • Uncertainty Quantification and Epistemic Humility. Current AI Scientists exhibit overconfidence, masking underlying uncertainties and alternative explanatory hypotheses. Principled uncertainty quantification requires architectures that explicitly model epistemic uncertainty (knowledge limitations) and aleatoric uncertainty (inherent stochasticity) throughout the research workflow. Promising approaches include Bayesian neural networks for confidence-calibrated predictions, ensemble methods maintaining diverse hypothesis sets [Yamada2025AIScientistV2], and conformal prediction providing distribution-free uncertainty intervals. Beyond technical mechanisms, systems must demonstrate epistemic humility—recognizing knowledge boundaries, expressing calibrated confidence, requesting additional experimentation when evidence is insufficient, and deferring to domain experts for decisions requiring contextual judgment. This shift from deterministic outputs to probabilistic reasoning with explicit uncertainty bounds is essential for trustworthy scientific automation.
@@
  • Cross-Domain Generalization through Modular Composition. Current AI Scientists achieve domain-specific success [Boiko2023Coscientist] but exhibit limited transferability to fields with informal methodologies or rapidly evolving instrumentation. Monolithic end-to-end architectures encode domain assumptions implicitly, preventing compositional reuse of learned capabilities. Cross-domain generalization requires transition toward modular, composable architectures based on capability factorization. This paradigm decomposes autonomous research into reusable functional modules—causal inference, symbolic regression, statistical testing, data visualization—each with formally specified input-output contracts. A meta-planning layer dynamically composes these modules into domain-adapted workflows, enabling procedural transfer through novel module assemblies rather than domain-specific retraining. Standardization of module interfaces, analogous to software engineering’s API contracts, would facilitate community-driven development of specialized capabilities while preserving interoperability. This architectural shift from monolithic systems to composable toolkits represents a prerequisite for general-purpose scientific automation.
@@
  • Human-AI Collaboration and Ethical Governance. The frontier of impactful research automation lies not in full autonomy but in synergistic human-AI partnerships that leverage complementary strengths—human creativity, intuition, and contextual judgment combined with AI’s scalability, systematic exploration, and computational power. Frameworks exemplified by freephdlabor [Li2025FreephdLabor] demonstrate interactive research automation where human strategists guide exploration while AI executors perform systematic experimentation and analysis. Realizing this vision requires formal protocols for role allocation, mixed-initiative planning, and explanation generation enabling meaningful human oversight. Concurrently, ethical governance frameworks become imperative as AI-generated research enters the scientific record. Essential components include machine-readable author contribution statements distinguishing human intellectual input from AI execution [resnik2025ethics], cryptographically verifiable audit trails documenting decision provenance, and risk-gated execution mechanisms requiring human approval for high-stakes domains [ateriya2025exploring]. Establishing community-wide standards for AI authorship attribution, intellectual property rights, and accountability mechanisms represents both a technical challenge—requiring standardized metadata schemas and verification protocols—and a societal imperative for maintaining scientific integrity and public trust in AI-augmented research.

### v5

@@ 4.4 Physics and Engineering
    SR-Scientist [64] employs agentic workflows for symbolic regression, autonomously discovering fundamental equations from observational data.
@@
    AI Feynman [58] introduces physics-constrained symbolic search (dimensional consistency), rediscovering classical laws from raw data and inspiring modern symbolic reasoning architectures.
@@
    Quantum-Agent-SDL [6] combines reinforcement learning with LLM-based hypothesis refinement for self-optimized qubit calibration and error correction in quantum computing.
@@ 4.5 Meta-Science and Social Science
    SciAgents [14] employs multi-agent collaboration and dynamic graph reasoning to traverse publication/author networks, identifying knowledge gaps, interdisciplinary connections, and predicting research trends.
@@
    AI for Social Science (AI4SS) [65] outlines roadmaps for AI-driven large-scale social modeling, policy simulation, and analysis of innovation diffusion and collaboration patterns.
@@
    Ethical Governance Frameworks [45, 41] address societal implications of AI-generated science, proposing frameworks for authorship management, credit attribution, and accountability maintenance.
@@ 5 Open Problems and Future Directions
  • Reproducibility-by-Design and Verifiable Science. Multi-stage autonomous systems amplify reproducibility challenges through compositional complexity, where parameter variations in early stages cascade into divergent outcomes. Achieving verifiable science requires architectural integration of three components: (1) Environmental Determinism—containerization of computational dependencies, cryptographic hashing of data artifacts, and version control of model checkpoints ensure bitwise reproducibility; (2) Fine-Grained Provenance—cryptographic linking of manuscript claims to specific code commits, data versions, and model parameters enables claim-level verification [17]; and (3) Automated Verification—AI-driven auditor agents performing logical consistency checks, detecting claim-evidence misalignments, and validating statistical inferences before publication [51]. These mechanisms must transition from post-hoc validation to intrinsic architectural properties.
@@
  • Uncertainty Quantification and Epistemic Humility. Current AI Scientists exhibit overconfidence, masking underlying uncertainties and alternative explanatory hypotheses. Principled uncertainty quantification requires architectures that explicitly model epistemic uncertainty (knowledge limitations) and aleatoric uncertainty (inherent stochasticity) throughout the research workflow. Promising approaches include Bayesian neural networks for confidence-calibrated predictions, ensemble methods maintaining diverse hypothesis sets [66], and conformal prediction providing distribution-free uncertainty intervals. Beyond technical mechanisms, systems must demonstrate epistemic humility—recognizing knowledge boundaries, expressing calibrated confidence, requesting additional experimentation when evidence is insufficient, and deferring to domain experts for decisions requiring contextual judgment. This shift from deterministic outputs to probabilistic reasoning with explicit uncertainty bounds is essential for trustworthy scientific automation.
@@
  • Cross-Domain Generalization through Modular Composition. Current AI Scientists achieve domain-specific success [5] but exhibit limited transferability to fields with informal methodologies or rapidly evolving instrumentation. Monolithic end-to-end architectures encode domain assumptions implicitly, preventing compositional reuse of learned capabilities. Cross-domain generalization requires transition toward modular, composable architectures based on capability factorization. This paradigm decomposes autonomous research into reusable functional modules—causal inference, symbolic regression, statistical testing, data visualization—each with formally specified input-output contracts. A meta-planning layer dynamically composes these modules into domain-adapted workflows, enabling procedural transfer through novel module assemblies rather than domain-specific retraining. Standardization of module interfaces, analogous to software engineering’s API contracts, would facilitate community-driven development of specialized capabilities while preserving interoperability. This architectural shift from monolithic systems to composable toolkits represents a prerequisite for general-purpose scientific automation.
@@
  • Human-AI Collaboration and Ethical Governance. The frontier of impactful research automation lies not in full autonomy but in synergistic human-AI partnerships that leverage complementary strengths—human creativity, intuition, and contextual judgment combined with AI’s scalability, systematic exploration, and computational power. Frameworks exemplified by freephdlabor [29] demonstrate interactive research automation where human strategists guide exploration while AI executors perform systematic experimentation and analysis. Realizing this vision requires formal protocols for role allocation, mixed-initiative planning, and explanation generation enabling meaningful human oversight. Concurrently, ethical governance frameworks become imperative as AI-generated research enters the scientific record. Essential components include machine-readable author contribution statements distinguishing human intellectual input from AI execution [45], cryptographically verifiable audit trails documenting decision provenance, and risk-gated execution mechanisms requiring human approval for high-stakes domains [4]. Establishing community-wide standards for AI authorship attribution, intellectual property rights, and accountability mechanisms represents both a technical challenge—requiring standardized metadata schemas and verification protocols—and a societal imperative for maintaining scientific integrity and public trust in AI-augmented research.